OpenStudy (anonymous):

How long does it take for radiation from a cesium-133 atom to complete 1 petacycle? in miliseconds

OpenStudy (aum):

Frequency of Cesium-133 is 9.192631770 x 10^9 Hertz or cycles / second. 1 petacycle = 10^15 cycles. Time taken in seconds = \(\Large \frac {10^{15}}{9.192631770 \times 10^9} = \frac{10^{6}}{9.192631770} = 108,783\) seconds To convert to milliseconds multiply by 1000 or 10^3: \(\large 1.08783 \times 10^8\) milliseconds.